if the average person drinks 8, (8oz) glasses of water per day, a person who drinks 12.8 oz of water after a morning exercise session has consumed what fraction of the daily average?
Answer by checkley77(12844)

12.8/8*8=12.8/64=.2=20% of the daily average.